OpenStudy (anonymous):

In a polygon, is it possible for each interior angle to be equal to 125?

OpenStudy (anonymous):

180(n-2) where n is the number of sides.

OpenStudy (anonymous):

If I do (n-2)180/n, I end up getting a decimal.

OpenStudy (anonymous):

\[{180 \cdot (n-2) \over n}=125\]\[125 n = 180n-360\]\[55n = 360\]\[n = {360 \over 55} = {72 \over 11}\] If you get a decimal, it means that it isn't possible.